# Neodymium ions as charge reservoir in NdNiO2: from lack of long range order to electron-doping-induced antiferromagnetism

### v1

@@ IV.1 Ferromagnetic and antiferromagnetic phases
Just as in the FM case, the stability of the AFM phase is due to relatively large Hubbard $U$. In this case, the on-site Coulomb interaction also splits the nickel bands into an upper and a lower band [Fig. 3(b, d, f)], opening a substantial gap, equal to $2U\rho_{1}$. The neodymium DOS remains in that gap [Fig. 3(d, f)]. However, for the same realistic parameter values as in the FM case ($\epsilon=2$ eV, $n=1$), in the AFM case there is almost no overlap between the lower nickel band DOS and the neodymium DOS. Consequently, for $n\gtrsim 1$ all charge occupies the lower nickel band, leading to half-filling on nickel and stabilising the AFM solution.
@@
Varying $\epsilon$, $n=1$. Changing the charge transfer energy $\epsilon$ [the vertical axis of Fig. 1] amounts to shifting the neodymium DOS upwards or downwards in energy [see Fig. 3, panels (e) and (f)]. In the FM case, for ϵ<italic-$\epsilon<$ 2eV, i.e. lower than the realistic value of $\epsilon=2$ eV, the overlap between the lower nickel Stoner band DOS and the neodymium DOS is increased; on the other hand, for $\epsilon>2$ eV the overlap is decreased. This decrease in the overlap for increasing $\epsilon$ leads to the lower nickel band eventually becoming fully occupied and the FM state becoming energetically less favorable (see Fig. 1) and ultimately unstable for large enough $\epsilon$ (unshown).
@@ Acknowledgements
We thank Karsten Held, Marcin Wysokiński and Adnrzej M. Oleś for stimulating discussions. We kindly acknowledge support from the National Science Centre (NCN, Poland) under Project No. 2021/43/B/ST3/02166. The work is supported by the Foundation for Polish Science through the IRA Programme co-financed by EU within SG OP Programme. A.M.O. is grateful for support via the Alexander von Humboldt Foundation Fellowship (Humboldt-Forschungspreis).

### v2

@@ IV.1 Ferromagnetic and antiferromagnetic phases
Just as in the FM case, the stability of the AFM phase is due to relatively large Hubbard $U$. In this case, the on-site Coulomb interaction also splits the nickel bands into an upper and a lower band [Fig. 6(b, d, f)], opening a substantial gap, equal to $2U\rho_{1}$. The neodymium DOS remains in that gap [Fig. 6(d, f)]. However, for the same realistic parameter values as in the FM case ($\epsilon=2$ eV, $n=1$), in the AFM case there is almost no overlap between the lower nickel band DOS and the neodymium DOS. Consequently, for $n\gtrsim 1$ all charge occupies the lower nickel band, leading to half-filling on nickel and stabilizing the AFM solution.
@@
Varying $\epsilon$, $n=1$. Changing the charge transfer energy $\epsilon$ [the vertical axis of Fig. 2] amounts to shifting the neodymium DOS upwards or downwards in energy [see Fig. 6, panels (e) and (f)]. In the FM case, for $\epsilon<$ 2eV, i.e. lower than the realistic value of $\epsilon=2$ eV, the overlap between the lower nickel Stoner band DOS and the neodymium DOS is increased; on the other hand, for $\epsilon>2$ eV the overlap is decreased. This decrease in the overlap for increasing $\epsilon$ leads to the lower nickel band eventually becoming fully occupied and the FM state becoming energetically less favorable (see Fig. 2) and ultimately unstable for large enough $\epsilon$ (unshown).
@@ V Conclusions
It is also interesting to note that the more complex Hartree-Fock calculations (using the Broyden mixing method) may also support the onset of stripes. This shows that stripes can become stable already on a static mean-field level and the more advanced treatment of correlations on the DMFT level is not needed to observe stripes in these systems [54]. The onset of stripes occurs for relatively low charge-transfer energies and for particular fillings. While in general stripes can always be expected in the doped Hubbard-like system, it is interesting to note that here they actually arise due to the self-doping effect. Note, however, that for a realistic case of the charge transfer energy $\epsilon=2$ eV, i.e. being not too low, the self-doping effect is small and according to the mean-field calculations of this work the stripes cannot become stable neither for pristine nor for electron-doped NdNiO2.
@@ Acknowledgements
We thank Karsten Held, Marcin Wysokiński and Andrzej M. Oleś for stimulating discussions. We kindly acknowledge support from the National Science Centre (NCN, Poland) under Project No. 2021/43/B/ST3/02166 and 2024/55/B/ST3/03144. The work is supported by the Foundation for Polish Science through the IRA Programme co-financed by EU within SG OP Programme.